Prove that m^r can be expressed as sum of m consecutive odd integers. Find the first odd integer in case of 8^5. m and r are natural numbers where r>1.

Grade:12

2 Answers

mycroft holmes
7 years ago
mr = (2a-1)+....+(2b-1) = b2- (a-1)= [b-(a-1)][b+a-1] = mr
 
We are given that b-(a-1) = m which means b+(a-1) = mr-1
 
So choose a such that a = m(mr-2-1)/2 + 1 and b = m(mr-2+1)/2
 
In case of 85 it is 2045
 
GAGANDEEP SINGH
7 years ago
Yeah solved it a little later. Thanks for sparing for valuable time on this. But your answer to 8^5 is wrong. This was a question in Sequences and Series.